# Flavored QCD axion and Modular invariance

### v2

@@ II Minimal set-up
Under the modular group $SL(2,\mathbb{Z})$ and the gauged $U(1)_{X}$, for the action (2) to be invariant we consider a low-energy Kähler potential $K$, superpotential $W$, and gauge kinetic function $f_{ab}$:
$K$ $=$ $-M^{2}_{P}\ln\Big\{(-i\tau+i\bar{\tau})^{h}\big(S+\bar{S}-\frac{h}{4\pi^{2}}\ln(-i\tau+i\bar{\tau})\big)\Big(U_{X}+\bar{U}_{X}-\frac{\delta^{\rm GS}_{X}}{16\pi^{2}}V_{X}\Big)^{h}\Big\}$
@@
$f_{ab}$ $=$ $\delta_{ab}(S+U_{X})\,$, (4)
with $h=3$ and the $U(1)_{X}$ charge $|X|=1$ determined by the transformation of $\varphi_{X}$, where k-k is the modular weight of $\varphi$, $Z_{X}$ is the normalization factor, $S$ denotes the axio-dilaton, $\tau$ represents the Kähler modulus, $U_{X}$ corresponds to the complex structure modulus, and the modular- and gauge-invariant superpotential $W(S,U_{X},\tau)$ is given by Eq.(33). The dots in Eq.(4) denote the contributions of non-renomalizable terms scaled by an UV cutoff $M_{P}$. The $U(1)_{X}$ charged matter fields $\varphi_{X}$ and complex structure modulus $U_{X}$ and the vector superfield $V_{X}$ of the gauged $U(1)_{X}$ containing the gauge field $A^{\mu}_{X}$ participate in the 4D Green-Schwarz (GS) mechanism Green:1984sg . The axio-dilaton $S$, the $U(1)_{X}$ charged modulus $U_{X}$, and the $U(1)_{X}$ charged scalar field $\varphi_{X}$ can be decomposed as  [^3] The superpotential $W(S,U_{X},\tau)$ of Eq.(33) explicitly breaks the $S$-shift symmetry.
$S=\frac{1}{g^{2}_{s}}+i\theta_{s}\,,\qquad U_{X}=\sigma+i\theta_{X}\,,\qquad\varphi_{X}\big|_{\theta=\bar{\theta}=0}=\frac{1}{\sqrt{2}}e^{i\frac{A_{X}}{v_{X}}}(v_{X}+h_{X})\,$, (5)
where  [^4] Assuming ${\rm Re}(U_{X})\gg{\rm Re}(S)$, the dilaton $S$ couples to all SM gauge groups with equal strength at tree level, as set by the string scale. However, this universality is broken by quantum effects, including threshold corrections and renormalization group running. $\sigma=1/g^{2}_{X}$ with $g_{X}$ being the 4D gauge coupling of $U(1)_{X}$, and $A_{X}$, $v_{X}$, and $h_{X}$ are the Nambu-Goldstone (NG) mode, VEV, and Higgs boson of scalar components, respectively. The GS parameter $\delta^{\rm GS}_{X}$ characterizes the coupling of the anomalous gauge boson to the closed string axion $\theta_{X}$.
@@ II.1 $SL(2,\mathbb{Z})$ modular anomaly cancellation
$\varphi_{i}\rightarrow(c\tau+d)^{-k_{i}}\varphi_{i}\,,\qquad S\rightarrow S-\frac{1}{4\pi^{2}}\ln(c\tau+d)^{h}\,$, (6)
@@
This normalization leads to $W=\frac{1}{3}\hat{Y}(\tau)\hat{\varphi}_{i}\hat{\varphi}_{j}\hat{\varphi}_{k}$ with the normalized modular form given  [^5] The physical modular forms can be expressed in a good approximation for $M_{P}\gg\langle\varphi_{X}\rangle,\langle\varphi\rangle$ and $V_{X}=0$ as $\hat{Y}(\tau)\simeq Y(\tau)(-i\tau+i\bar{\tau})^{(k_{i}+k_{j}+k_{k}-h)/2}(U_{X}+\bar{U}_{X})^{-\frac{h}{2}}(S+\bar{S}-\frac{h}{4\pi^{2}}\ln(-i\tau+i\bar{\tau}))^{-\frac{1}{2}}$. Without loss of generality, the term $(U_{X}+\bar{U}_{X})^{-\frac{h}{2}}(S+\bar{S}-\frac{h}{4\pi^{2}}\ln(-i\tau+i\bar{\tau}))^{-\frac{1}{2}}$ can be absorbed to Yukawa coefficients and normalized to one, leading to $\hat{Y}(\tau)\simeq Y(\tau)(-i\tau+i\bar{\tau})^{(k_{i}+k_{j}+k_{k}-h)/2}$. by $\hat{Y}(\tau)=e^{K/2M^{2}_{P}}Y(\tau)(-i\tau+i\bar{\tau})^{(k_{i}+k_{j}+k_{k})/2}$ within the framework of supergravity. Under the modular transformations of Eqs.(1) and (6), the normalized fields and modular forms transform as
@@
where $F^{i}=e^{K/2M^{2}_{P}}K^{i\bar{j}}\bar{D}_{\bar{j}}\bar{W}$, $\sigma^{\mu}=(1,\sigma^{k})$ and $\bar{\sigma}^{\mu}=(1,-\sigma^{k})$ with $\sigma^{k}$ the Pauli matrices, $\Gamma^{i}_{jk}=K^{i\bar{\ell}}\partial_{j}K_{k\bar{\ell}}$ is the modular connection (Kähler Christoffel symbol), $K^{i\bar{j}}=(\partial_{i}\partial_{\bar{j}}K)^{-1}$ is the inverse Kähler metric, the Kähler-covariant derivative of the superpotential is $D_{i}W=W_{i}+\frac{K_{i}}{M^{2}_{P}}W$, and the fully covariant second derivative  [^6] The covariant derivative ${\cal D}_{i}$ acting on an object $V_{j}$ with Kähler weight $(p,q)$, meaning it transforms as $V_{j}\rightarrow e^{-(pg+q\bar{g})}V_{j}$ under the Kähler transformation Eq.(3), is given by ${\cal D}_{i}V_{j}=\partial_{i}V_{j}-\Gamma^{k}_{ij}V_{k}+\frac{p}{M^{2}_{P}}K_{i}V_{j}$. Here, under the Kähler transformation Eq.(3), the object $D_{j}W$ transforms as $D_{j}W\rightarrow e^{-g(\Phi)}D_{j}W$. is ${\cal D}_{i}D_{j}W=W_{ij}+\frac{K_{ij}}{M^{2}_{P}}W+\frac{K_{i}}{M^{2}_{P}}D_{j}W+\frac{K_{j}}{M^{2}_{P}}D_{i}W-\frac{K_{i}K_{j}}{M^{4}_{P}}W-\Gamma^{k}_{ij}D_{k}W$. The spacetime covariant derivatives acting on SM fermions and gauginos are given by
@@
$\Psi\rightarrow e^{-q_{K}\frac{g-\bar{g}}{2}}\Psi\,,\qquad\bar{\Psi}\rightarrow e^{-q_{K}\frac{g-\bar{g}}{2}}\bar{\Psi}\qquad\text{with}~\Psi=\hat{\psi},\hat{\lambda}\,$. (13)
@@
First, the anomaly generated by the chiral rotation of gauginos, which exactly matches that from the Kähler transformation  [^9] Under the Kähler transformation, the modular anomaly manifests via the variation of the action $\delta S=-\tilde{c}\frac{1}{4}\int d^{4}xd^{2}\theta{\cal W}^{\alpha}{\cal W}_{\alpha}g(\tau)+{\rm h.c.}$Derendinger:1991hq ., takes the form
$-\frac{C}{32\pi^{2}}\Big\{-\big(g(\tau)+g(\bar{\tau})\big)Q^{\mu\nu}Q_{\mu\nu}+i\big(g(\tau)-g(\bar{\tau})\big)Q^{\mu\nu}\tilde{Q}_{\mu\nu}\Big\}\,$, (17)
where the first term in the brackets corresponds to the gauge boson kinetic term, while the second, CP-odd term involves the dual field strength $\tilde{Q}_{\mu\nu}=\frac{1}{2}\epsilon_{\mu\nu\rho\sigma}Q^{\rho\sigma}$. The gauge field strengths $Q$ are given by $\{G,W,Y,F_{X}\}$ for $SU(3)_{C}$, $SU(2)_{L}$, $U(1)_{Y}$, and $U(1)_{X}$, respectively. In Eq.(17), the coefficients take values $C=3,2,1,1$ for gluino, wino, bino, and $U(1)_{X}$ gaugino, respectively. The anomalies are cancelled via the transformation of the modulus $S$ in Eq.(6), which enters the one-loop corrected gauge kinetic function,
$f^{\rm 1-loop}_{ab}(\Phi)\supset\delta_{ab}\big\{S-\frac{1}{4\pi^{2}}\ln(c\tau+d)^{h}\big\}\,$, (18)
as discussed in Refs.Ferrara:1989bc ; Ahn:2023iqa ; Derendinger:1991hq ; Ibanez:1992hc . The gaugino masses $M_{\hat{\lambda}}$ transform according to $M_{\hat{\lambda}}\rightarrow\Big(\frac{c\tau+d}{c\bar{\tau}+d}\Big)^{\frac{h}{2}}M_{\hat{\lambda}}$ due to Eq.(15). Since gauginos do not mix with SM fermions, their contribution to the modular anomalies cancels through the one-loop corrected gauge kinetic function Ahn:2023iqa , yielding
@@
where $\vartheta_{\rm QCD}$ comes from $\theta_{s}$ and $\theta_{X}$ of Eq.(5), and $A_{C}$ denotes the sum of the modular weights of the quark fields, under the assumption  [^10] We assume that all symmetry-breaking scalars have modular weight zero; otherwise their VEVs would have to be zero in order to remain invariant under the modular transformation, see Eqs.(6) or (9) and Eq.(47). $k_{H_{u(d)}}=0$ for the Higgs doublets $H_{u(d)}$.
@@ II.2 Gauged $U(1)_{X}$ anomaly-free and global $U(1)_{X}$
$K_{U_{X}\bar{U}_{X}}\Big(\partial^{\mu}\theta_{X}-\frac{\delta^{\rm GS}_{X}}{16\pi^{2}}A^{\mu}_{X}\Big)^{2}-\frac{1}{4g^{2}_{X}}F^{\mu\nu}_{X}F_{X\mu\nu}+g_{X}\xi^{\rm FI}_{X}D_{X}-D_{X}g_{X}X|\varphi_{X}|^{2}$
$+\frac{1}{2}D^{2}_{X}+|D_{\mu}\varphi_{X}|^{2}+\theta_{X}{\rm Tr}(Q^{\mu\nu}\tilde{Q}_{\mu\nu})+\frac{A_{X}}{f_{X}}\frac{\delta^{Q}_{X}}{16\pi^{2}}{\rm Tr}(Q^{\mu\nu}\tilde{Q}_{\mu\nu})\,$, (29)
where $Q=G,W,Y,F_{X}$ denote the gauge field strengths for $SU(3)_{C}$, $SU(2)_{L}$, $U(1)_{Y}$, and $U(1)_{X}$, respectively, with gauge couplings absorbed into their definitions. The first, third, fourth, and sixth terms result from expanding the Kähler potential Eq.(4), and the second, fifth, seventh, and eighth terms result from the gauge interaction term in Eq.(2). $F^{\mu\nu}_{X}=\partial^{\mu}A^{\nu}_{X}-\partial^{\nu}A^{\mu}_{X}$ is the $U(1)_{X}$ gauge field strength. The $U(1)_{X}$ gauge covariant derivative $D^{\mu}\varphi_{X}=\partial^{\mu}\varphi_{X}-iXA^{\mu}_{X}\varphi_{X}$ governs the coupling of the scalar component of $\varphi_{X}$ to the $U(1)_{X}$ gauge boson, where the gauge coupling $g_{X}$ is absorbed into $A^{\mu}_{X}$. The first and seventh terms together, and the sixth and eighth terms in Eq.(29), are gauge invariant under the anomalous $U(1)_{X}$ transformations of Eqs.(27) and (28). Under the gauge transformations of Eqs.(27) and (28), it requires
$\delta^{\rm GS}_{X}=\alpha^{Q}_{X}\,\delta^{Q}_{X}\,,\qquad\partial_{\mu}J^{\mu}_{X}=\frac{\delta^{\rm GS}_{X}}{16\pi^{2}}{\rm Tr}(Q^{\mu\nu}\tilde{Q}_{\mu\nu})=-\partial_{\mu}J^{\mu}_{\theta}\,$, (30)
@@
where the trace is over all fermions $\psi$ carrying $U(1)_{X}$ charges, $T_{SU(3)}$, $T_{SU(2)}$ are gauge group generators, $Y$ is the hypercharge operator, and $X_{\psi}$ denotes the $U(1)_{X}$ charge of $\psi$. Here, for convenience, $\delta^{Y}_{X}$ and $\delta^{F_{X}}_{X}$ are defined as above for hypercharge and $U(1)_{X}$, respectively. The Fayet-Iliopoulos (FI) term ${\cal L}^{\rm FI}_{X}=\xi^{\rm FI}_{X}\int d^{2}\theta d^{2}\bar{\theta}V_{X}=\xi^{\rm FI}_{X}g_{X}D_{X}$ with $D_{X}=g_{X}(-\xi^{\rm FI}_{X}+X|\varphi_{X}|^{2})$ leads to $D$-term potential $V_{D}=\frac{g^{2}_{X}}{2}(-\xi^{\rm FI}_{X}+X|\varphi_{X}|^{2})^{2}$ for the anomalous $U(1)_{X}$. Here $\xi^{\rm FI}_{X}$ is the FI factor $\xi^{\rm FI}_{X}=\frac{\partial K}{\partial V_{X}}|_{V_{X}=0,\sigma=\sigma_{0}}\Delta\sigma$ produced by expanding the Kähler potential Eq.(4) in components linear in $V_{X}$:

### v3

@@ II Minimal set-up
Under the modular group $SL(2,\mathbb{Z})$ and the gauged $U(1)_{X}$, for the action (2) to be invariant we consider a low-energy Kähler potential $K$, superpotential $W$, and gauge kinetic function $f_{Qab}$ (here $Q$ labels the gauge group factor):
$K$ $=$ $-M^{2}_{P}\ln\Big\{(-i\tau+i\bar{\tau})^{h}\big(S+\bar{S}-\frac{h}{16\pi^{2}}\ln(-i\tau+i\bar{\tau})\big)\Big(U_{X}+\bar{U}_{X}-\frac{\delta^{\rm GS}_{X}}{16\pi^{2}}V_{X}\Big)^{h}\Big\}$
@@
$f_{Qab}$ $=$ $\delta_{ab}(\kappa_{Q}\,S+\tilde{\kappa}_{Q}\,U_{X})\,$, (4)
with $h=3$ and the $U(1)_{X}$ charge $|X|=1$ determined by the transformation of $\varphi_{X}$, where k-k is the modular weight of $\varphi$, $Z_{X}$ is the normalization factor, $S$ denotes the dilaton, $\tau$ represents the Kähler modulus, $U_{X}$ corresponds to the complex structure modulus, and the modular- and gauge-invariant superpotential $W(S,U_{X},\tau)$ is given by Eq.(34). The dots in Eq.(4) denote the contributions of non-renomalizable terms scaled by an UV cutoff $M_{P}$. The $U(1)_{X}$ charged matter fields $\varphi_{X}$ and complex structure modulus $U_{X}$ and the vector superfield $V_{X}$ of the gauged $U(1)_{X}$ containing the gauge field $A^{\mu}_{X}$ participate in the 4D Green-Schwarz (GS) mechanism Green:1984sg . The dilaton $S$, the $U(1)_{X}$ charged modulus $U_{X}$, and the $U(1)_{X}$ charged scalar field $\varphi_{X}$ can be decomposed as
$S=\frac{1}{g^{2}_{s}}\,,\qquad U_{X}=\sigma+i\theta_{X}\,,\qquad\varphi_{X}\big|_{\theta=\bar{\theta}=0}=\frac{1}{\sqrt{2}}e^{i\frac{A_{X}}{v_{X}}}(v_{X}+h_{X})\,$, (5)
where $\sigma=1/g^{2}_{X}$ with $g_{X}$ being the 4D gauge coupling of $U(1)_{X}$, and $A_{X}$, $v_{X}$, and $h_{X}$ are the Nambu-Goldstone (NG) mode, VEV, and Higgs boson of scalar components, respectively. The GS parameter $\delta^{\rm GS}_{X}$ characterizes the coupling of the anomalous gauge boson to the closed string axion $\theta_{X}$.
@@ II.1 $SL(2,\mathbb{Z})$ modular anomaly cancellation
$\varphi_{i}\rightarrow(c\tau+d)^{-k_{i}}\varphi_{i}\,,\qquad S\rightarrow S-\frac{1}{16\pi^{2}}\ln(c\tau+d)^{h}\,$, (6)
@@
This normalization leads to $W=\frac{1}{3}\hat{Y}(\tau)\hat{\varphi}_{i}\hat{\varphi}_{j}\hat{\varphi}_{k}$ with the normalized modular form given  [^3] The physical modular forms can be expressed in a good approximation for $M_{P}\gg\langle\varphi_{X}\rangle,\langle\varphi\rangle$ and $V_{X}=0$ as $\hat{Y}(\tau)\simeq Y(\tau)(-i\tau+i\bar{\tau})^{(k_{i}+k_{j}+k_{k}-h)/2}(U_{X}+\bar{U}_{X})^{-\frac{h}{2}}(S+\bar{S}-\frac{h}{16\pi^{2}}\ln(-i\tau+i\bar{\tau}))^{-\frac{1}{2}}$. Without loss of generality, the term $(U_{X}+\bar{U}_{X})^{-\frac{h}{2}}(S+\bar{S}-\frac{h}{16\pi^{2}}\ln(-i\tau+i\bar{\tau}))^{-\frac{1}{2}}$ can be absorbed to Yukawa coefficients and normalized to one, leading to $\hat{Y}(\tau)\simeq Y(\tau)(-i\tau+i\bar{\tau})^{(k_{i}+k_{j}+k_{k}-h)/2}$. by $\hat{Y}(\tau)=e^{K/2M^{2}_{P}}Y(\tau)(-i\tau+i\bar{\tau})^{(k_{i}+k_{j}+k_{k})/2}$ within the framework of supergravity. Under the modular transformations of Eqs.(1) and (6), the normalized fields and modular forms transform as
@@
where $F^{i}=-e^{K/2M^{2}_{P}}K^{i\bar{j}}\bar{D}_{\bar{j}}\bar{W}$, $\sigma^{\mu}=(1,\sigma^{k})$ and $\bar{\sigma}^{\mu}=(1,-\sigma^{k})$ with $\sigma^{k}$ the Pauli matrices, $\Gamma^{i}_{jk}=K^{i\bar{\ell}}\partial_{j}K_{k\bar{\ell}}$ is the modular connection (Kähler Christoffel symbol), $K^{i\bar{j}}=(\partial_{i}\partial_{\bar{j}}K)^{-1}$ is the inverse Kähler metric, the Kähler-covariant derivative of the superpotential is $D_{i}W=W_{i}+\frac{K_{i}}{M^{2}_{P}}W$, and the fully covariant second derivative  [^4] The covariant derivative ${\cal D}_{i}$ acting on an object $V_{j}$ with Kähler weight $(p,q)$, meaning it transforms as $V_{j}\rightarrow e^{-(pg+q\bar{g})}V_{j}$ under the Kähler transformation Eq.(3), is given by ${\cal D}_{i}V_{j}=\partial_{i}V_{j}-\Gamma^{k}_{ij}V_{k}+\frac{p}{M^{2}_{P}}K_{i}V_{j}$. Here, under the Kähler transformation Eq.(3), the object $D_{j}W$ transforms as $D_{j}W\rightarrow e^{-g(\Phi)}D_{j}W$. is ${\cal D}_{i}D_{j}W=W_{ij}+\frac{K_{ij}}{M^{2}_{P}}W+\frac{K_{i}}{M^{2}_{P}}D_{j}W+\frac{K_{j}}{M^{2}_{P}}D_{i}W-\frac{K_{i}K_{j}}{M^{4}_{P}}W-\Gamma^{k}_{ij}D_{k}W$. The spacetime covariant derivatives acting on SM fermions and gauginos are given by
@@
$\Psi\rightarrow e^{-q_{K}\frac{g-\bar{g}}{2}}\Psi\,,\qquad\bar{\Psi}\rightarrow e^{q_{K}\frac{g-\bar{g}}{2}}\bar{\Psi}\qquad\text{with}~\Psi=\hat{\psi},\hat{\lambda}\,$. (13)
@@
First, we consider how the $SL(2,\mathbb{Z})\times[G_{Q}]^{2}$ anomalies induced by the chiral rotation of gauginos in Eq.(15) are cancelled via the GS mechanism, see also Refs.Ferrara:1989bc ; Ahn:2023iqa ; Derendinger:1991hq ; Ibanez:1992hc . Under the dilaton transformation in Eq.(6), the gauge kinetic function in Eq.(4) transforms as $f_{Q}\rightarrow f_{Q}-\frac{\kappa_{Q}}{16\pi^{2}}\ln(c\tau+d)^{h}$. This induces the following variation of the action,
$\delta S_{S}=\int d^{4}x\Big(-\frac{\kappa_{Q}}{16\pi^{2}}\Big)\Big\{-\frac{1}{8}Q^{\mu\nu}Q_{\mu\nu}\big(g(\tau)+g(\bar{\tau})\big)+\frac{i}{8}Q^{\mu\nu}\tilde{Q}_{\mu\nu}\big(g(\tau)-g(\bar{\tau})\big)\Big\}\,$, (17)
which exactly matches the anomaly induced by the chiral rotation of the gauginos in Eq.(15). Here the first term in the brackets corresponds to the gauge boson kinetic term, while the second, CP-odd term involves the dual field strength $\tilde{Q}_{\mu\nu}=\frac{1}{2}\epsilon_{\mu\nu\rho\sigma}Q^{\rho\sigma}$. The gauge field strengths $Q$ are given by $\{G,W,Y,F_{X}\}$ for $SU(3)_{C}$, $SU(2)_{L}$, $U(1)_{Y}$, and $U(1)_{X}$, respectively. Meanwhile, the tree-level action including one-loop contribution from massless gauginos takes the form (see Ref.string_book ) $S\supset\int d^{4}xd^{2}\theta\frac{1}{4}{\cal W}^{\alpha}{\cal W}_{\alpha}\big(f_{Q}+\frac{C(G_{Q})}{16\pi^{2}}\frac{K}{M^{2}_{P}}\big)+{\rm h.c.}$, where $C(G_{Q})$ is the quadratic Casimir coefficient of each gauge group $G_{Q}$, taking values $3,2,0$ for $SU(3)$, $SU(2)$, and $U(1)$, respectively. Then, under the Kähler transformation  [^7] Under the Kähler transformation, the modular anomaly manifests via the variation of the action $\delta S=\tilde{c}\frac{1}{4}\int d^{4}xd^{2}\theta{\cal W}^{\alpha}{\cal W}_{\alpha}g(\tau)+{\rm h.c.}$Derendinger:1991hq . in Eq.(3) the action varies as
$\delta S_{K}=\int d^{4}x\frac{C(G_{Q})}{16\pi^{2}}\Big\{-\frac{1}{8}Q^{\mu\nu}Q_{\mu\nu}\big(g(\tau)+g(\bar{\tau})\big)+\frac{i}{8}Q^{\mu\nu}\tilde{Q}_{\mu\nu}\big(g(\tau)-g(\bar{\tau})\big)\Big\}\,$, (18)
Therefore the total variation of the action is $\delta S=\delta S_{S}+\delta S_{K}\propto C(G_{Q})-\kappa_{Q}$. Hence the anomaly vanishes when $C(G_{Q})=\kappa_{Q}$, which corresponds to the GS universality condition, see also Ref.Ibanez:1992hc . Therefore the modular anomalies are cancelled consistently when the coefficient $\kappa_{Q}$ equals the quadratic Casimir coefficient of each gauge group.
The gaugino masses $M_{\hat{\lambda}}$ transform according to $M_{\hat{\lambda}}\rightarrow\Big(\frac{c\tau+d}{c\bar{\tau}+d}\Big)^{\frac{h}{2}}M_{\hat{\lambda}}$ due to Eq.(15). Since gauginos do not mix with SM fermions, their contribution to the modular anomalies cancels through the one-loop corrected gauge kinetic function Ahn:2023iqa , yielding
@@
where $\vartheta_{\rm QCD}$ is the bare QCD vacuum angle, and $A_{C}$ denotes the sum of the modular weights of the quark fields, under the assumption  [^8] We assume that all symmetry-breaking scalars have modular weight zero; otherwise their VEVs would have to be zero in order to remain invariant under the modular transformation, see Eqs.(6) or (9) and Eq.(48). $k_{H_{u(d)}}=0$ for the Higgs doublets $H_{u(d)}$.
@@ II.2 Gauged $U(1)_{X}$ anomaly-free and global $U(1)_{X}$
$K_{U_{X}\bar{U}_{X}}\Big(\partial^{\mu}\theta_{X}-\frac{\delta^{\rm GS}_{X}}{16\pi^{2}}A^{\mu}_{X}\Big)^{2}-\frac{1}{4g^{2}_{Q}}Q^{\mu\nu}Q_{\mu\nu}+\tilde{g}_{X}\xi^{\rm FI}_{X}D_{X}-D_{X}\tilde{g}_{X}X|\varphi_{X}|^{2}$
$+\frac{1}{2}D^{2}_{X}+|D_{\mu}\varphi_{X}|^{2}-\tilde{\kappa}_{Q}\frac{\theta_{X}}{2}{\rm Tr}(Q^{\mu\nu}\tilde{Q}_{\mu\nu})-\tilde{\kappa}_{Q}\frac{A_{X}}{f_{X}}\frac{\delta^{Q}_{X}}{32\pi^{2}}{\rm Tr}(Q^{\mu\nu}\tilde{Q}_{\mu\nu})\,$, (29)
where $Q=G,W,Y,F_{X}$ denote the gauge field strengths for $SU(3)_{C}$, $SU(2)_{L}$, $U(1)_{Y}$, and $U(1)_{X}$, respectively, with gauge couplings absorbed into their definitions. From Eqs.(4) and (5) the gauge coupling constants at a scale $\Lambda$ are given by $1/g^{2}_{Q}=\kappa_{Q}/g^{2}_{s}+\tilde{\kappa}_{Q}/g^{2}_{X}$, where $g^{2}_{Q}\equiv g^{2}_{Q}(\Lambda)$, which suggests that the physical gauge couplings unify around the scale $\Lambda$. The coefficients $\kappa_{Q}$ and $\tilde{\kappa}_{Q}$ are fixed by GS universality conditions (see below Eq.(18) and footnote-8). For the $U(1)_{X}$ gauge group this gives
$\frac{1}{\tilde{g}^{2}_{X}}=\frac{\tilde{\kappa}_{X}}{g^{2}_{X}}\,$, (30)
since $C(G_{X})=0$ for an Abelian gauge group, consistent with the GS cancellation condition, see below Eq.(18). The first, third, fourth, and sixth terms result from expanding the Kähler potential Eq.(4), and the second, fifth, seventh, and eighth terms result from the gauge interaction term in Eq.(2). $F^{\mu\nu}_{X}=\partial^{\mu}A^{\nu}_{X}-\partial^{\nu}A^{\mu}_{X}$ is the $U(1)_{X}$ gauge field strength. The $U(1)_{X}$ gauge covariant derivative $D^{\mu}\varphi_{X}=\partial^{\mu}\varphi_{X}-iXA^{\mu}_{X}\varphi_{X}$ governs the coupling of the scalar component of $\varphi_{X}$ to the $U(1)_{X}$ gauge boson, where the gauge coupling $g_{X}$ is absorbed into $A^{\mu}_{X}$. The first and seventh terms together, and the sixth and eighth terms in Eq.(29), are gauge invariant under the anomalous $U(1)_{X}$ transformations of Eqs.(27) and (28). Under the gauge transformations of Eqs.(27) and (28), it requires  [^9] Under the $U(1)_{X}$ gauge transformations (see the text above Eq.(27)), the gauge kinetic function in Eq.(4) transforms as $f_{Q}\rightarrow f_{Q}+i\tilde{\kappa}_{Q}\frac{\delta^{\rm GS}_{X}}{16\pi^{2}}\Lambda_{X}$. This induces a variation the action $\delta S_{U_{X}}=\int d^{4}x\frac{1}{4}Q^{\mu\nu}\tilde{Q}_{\mu\nu}\big(\tilde{\kappa}_{Q}\frac{\delta^{\rm GS}_{X}}{16\pi^{2}}\xi_{X}\big)$. Meanwhile, under the chiral rotation of the SM fermions produces the anomaly contribution $\delta S_{U(1)_{X}}=\int d^{4}x\frac{1}{4}Q^{\mu\nu}\tilde{Q}_{\mu\nu}\big(-\frac{\delta^{Q}_{X}}{16\pi^{2}}\xi_{X}\big)$ from the path-integral measure. The total variation is therefore $\delta S=\int d^{4}x\big(-\frac{1}{4}\big)Q^{\mu\nu}\tilde{Q}_{\mu\nu}\frac{\xi_{X}}{16\pi^{2}}\big(\delta^{Q}_{X}-\tilde{\kappa}_{Q}\delta^{\rm GS}_{X}\big)$. Hence the anomaly cancels when $\delta^{Q}_{X}=\tilde{\kappa}_{Q}\delta^{\rm GS}_{X}$, which is exactly equivalent to the first equation in Eq.(31) for $\alpha^{Q}_{X}=1/\tilde{\kappa}_{Q}$.
$\delta^{\rm GS}_{X}=\alpha^{Q}_{X}\,\delta^{Q}_{X}\,,\qquad-\partial_{\mu}J^{\mu}_{X}=\tilde{\kappa}_{Q}\frac{\delta^{\rm GS}_{X}}{32\pi^{2}}{\rm Tr}(Q^{\mu\nu}\tilde{Q}_{\mu\nu})=\partial_{\mu}J^{\mu}_{\theta}\,$, (31)
@@
where the trace is over all fermions $\psi$ carrying $U(1)_{X}$ charges, $T_{SU(3)}$, $T_{SU(2)}$ are gauge group generators, $Y$ is the hypercharge operator, and $X_{\psi}$ denotes the $U(1)_{X}$ charge of $\psi$. Here, for convenience, $\delta^{Y}_{X}$ and $\delta^{F_{X}}_{X}$ are defined as above for hypercharge and $U(1)_{X}$, respectively. The Fayet-Iliopoulos (FI) term ${\cal L}^{\rm FI}_{X}=\xi^{\rm FI}_{X}\int d^{2}\theta d^{2}\bar{\theta}V_{X}=\xi^{\rm FI}_{X}\tilde{g}_{X}D_{X}$ with $D_{X}=\tilde{g}_{X}(-\xi^{\rm FI}_{X}+X|\varphi_{X}|^{2})$ leads to $D$-term potential $V_{D}=\frac{\tilde{g}^{2}_{X}}{2}(-\xi^{\rm FI}_{X}+X|\varphi_{X}|^{2})^{2}$ for the anomalous $U(1)_{X}$. Here $\xi^{\rm FI}_{X}$ is the FI factor $\xi^{\rm FI}_{X}=\frac{\partial K}{\partial V_{X}}|_{V_{X}=0,\sigma=\sigma_{0}}\Delta\sigma$ produced by expanding the Kähler potential Eq.(4) in components linear in $V_{X}$: